# Novel Lower Bounds on $M/G/k$ Scheduling

### v2

@@ 2.3. Basic adjoint relationship (BAR) and drift method
The well known Basic Adjoint Relationship (BAR) equation states that the stationary distribution $\pi$ of a continuous-time Markov chain $(Z_{t})_{t\geq 0}$ with instantaneous generator $G$ satisfies
@@
under suitable conditions on the Markov chain $Z$ and the function $g$ [6]. By carefully designing the test function $g$, one can solve the BAR equation (2.1) to obtain either exact expressions for the moments of stationary variables or (asymptotically tight) bounds on these moments, depending on the system. This method is commonly referred to as the drift method. See, e.g., [4, 25, 11, 18]. We use the drift method to obtain the exact expected total work of the ISQ-$k$ system, and exact results and bounds for further systems we introduce in Section 3. We use these exact results and bounds to lower bound mean response time in the $M/G/k$ under an arbitrary scheduling policy.
@@ 2.3.1. Choosing test functions
Recent applications of the drift method use simple exponential test functions of the form $e^{tQ}$, where $t$ is a general constant and $Q$ is the queue length, or $e^{tW}$, where $W$ is the total work in the system. For example, [1] and [2] use these test functions to prove heavy-traffic and steady-state approximations in both single-class and multi-class queueing systems. The use of exponential test functions is also known as the transform method, see, e.g., [19] and [20], which likewise uses test functions of the form $e^{tQ}$ or $e^{tW}$. Other recent works have focused on test functions of the form $Q^{2}$ or $W^{2}$, [10], which likewise depend on a constant rate of work completion.
However, these methods cannot be directly applied to queues with a variable work completion rate, as the test functions $e^{tQ}$ and $e^{tW}$ only yield useful information when the work completion rate is constant. Some attempts have been made to circumvent this problem by focusing on heavy traffic settings. For instance, [19] relies heavily on state-space collapse to ensure a constant work completion rate except when the system is near-empty. As the system is rarely empty in heavy traffic, this state-space collapse allows the same straightforward test functions to be used, and heavy traffic results to be obtained.
@@ 2.3.2. Sufficient conditions for BAR
Proposition 3 of [6] offers a proof of a set of sufficient conditions for the BAR equation (2.1) to hold. This proposition has been widely applied in many recent papers in queueing theory. Their result requires the state space to be discrete and certain regularity conditions to hold on the rate matrix of the Markov jump process and the test function $g$. For example, [11] and [18] invoke Proposition 3 of [6] by showing that their respective Markov chains have uniformly bounded transition rates. However, this result does not directly apply to our system because our system has a continuous state space.
Theorem 2 of [6] provides another set of sufficient conditions for the BAR equation (2.1) to hold. This result allows continuous state spaces but requires bounded-drift test functions. However, it is not directly applicable to our unbounded quadratic test function, which has unbounded drift. For test functions with unbounded drift, a standard approach in the existing literature is to truncate these functions, thereby producing a sequence of bounded test functions that approximate the original. See, e.g., [1] and [21].
@@ 3.1. Queueing model and notation
We say a job is relevant at some threshold $x$ if its remaining size is below $x$. This may occur if the job arrives into the system with an initial size $<x$, or if a job with an initial size $>x$ reaches a remaining size below $x$. We refer to the process of a job receiving service and lowering its remaining size as “aging”.
We use $W$ to denote the total work in the system, namely the sum of the remaining sizes of all jobs in the system. We use $W_{x}$ to denote the relevant work, the total remaining size of all relevant jobs in the system.
Let $\lambda_{x}:=\lambda F_{S}(x)$ denote the arrival rate of jobs which are relevant from the moment they enter the system. Let $S_{x}:=[S\mid S\leq x]$ denote the conditional job size distribution, truncated at a threshold $x$. The conditional load, $\rho_{x}$, at a threshold $x$ is the average rate at which works with initial size $<x$ arrives into the system is given by $\rho_{x}:=\lambda_{x}\mathbb{E}[S_{x}]=\lambda\mathbb{E}[S\mathbbm{1}_{\{S\leq x\}}]$.
We also define the capped job size distribution, $S_{\bar{x}}:=[\min\{S,x\}]$, which represents the amount of relevant work at some threshold $x$ that a random job will contribute over its time in the system. The capped load, $\rho_{\bar{x}}$, is similarly given by $\rho_{\bar{x}}=\lambda\mathbb{E}[S_{\bar{x}}]=\lambda\mathbb{E}[\min\{S,x\}]$.
@@ 3.2. Increasing speed queue
3.3. Recycling: auxiliary ISQ systems
@@
Recall that we call a job relevant if its remaining size is less than some threshold $x$. There are two ways for a job to be relevant: it either enters the system with size less than $x$, or it ages down from a size larger than $x$ to a size less than $x$. We refer to the latter as recycling, and we call such jobs recycled.
@@
We define two auxiliary systems to help us lower bound the response time of the $M/G/k$ system: the separate-recycling ISQ-$k$ and the arbitrary-recycling ISQ-$k$.
@@
The separate-recycling ISQ-$k$ system (Sep-ISQ-$k$) consists of two subsystems, the truncated-ISQ-$k$ subsystem and the $M/G/\infty$ subsystem. Jobs arrive to the overall Sep-ISQ-$k$ system with the same arrival process as the $M/G/k$, jobs arriving into the system according to a Poisson process with rate $\lambda$ and with job size distribution $S$. Jobs with initial sizes $\leq x$ are routed to an ISQ-$k$ subsystem, which we call the truncated-ISQ-$k$ subsystem. Specifically, the truncated-ISQ-$k$ system is an ISQ-$k$ system characterized by an arrival rate $\lambda_{x}=\lambda F_{S}(x)$, and job size distribution $S_{x}\sim[S\mid S\leq x]$.
@@
On the other hand, jobs with initial sizes $>x$ are routed to the separate $M/G/\infty$ subsystem, each server operating at a speed of $1/k$. That is, jobs which are not relevant on arrival are served at a separate subsystem.
@@
The arbitrary-recycling-ISQ-$k$ system (AR-ISQ-$k$) is an ISQ-$k$ system with two arrival streams, a Poisson arrival process with a rate of $\lambda F_{S}(x)$ and job size distribution $S_{x}\sim[S\mid S\leq x]$, and a general arrival process with rate $\lambda(1-F_{S}(x))$ and job sizes exactly $x$. The general arrival process is governed by some general Markovian process for which we only require stability and some general hidden state. The idea behind this general arrival process is that it models the moments at which jobs recycle in the M/G/$k$ system under a general Markovian scheduling policy. We refer to the Poisson arrivals as the truncated stream and the general arrival process as the recycling stream.
@@
We will use the triplet $(w,i,a)$ to denote a state of the arbitrary recycling system where $w$ and $i$ are defined similarly to the states of the full ISQ-$k$ system and $a$ denotes a separate Markovian state from the recycling stream’s state space, which incorporates additional information about the recycling stream. As before, $A$ denotes the corresponding stationary random variable.
@@
In Theorem 6.2, we show that the mean relevant work in the Sep-ISQ-$k$ system lower bounds the mean relevant work in the $M/G/k$ system under arbitrary scheduling policy, for any arrival rate $\lambda$, job size distribution $S$, and relevancy cutoff $x$. We also show that the minimum mean relevant work in the AR-ISQ-$k$ system over all recycling streams lower bounds the mean relevant work in the $M/G/k$ system under an arbitrary scheduling policy, see Theorem 6.3.
@@ 4. Main Results
In this paper, we give the first nontrivial lower bounds on the expected relevant work of the $M/G/k$ system under an arbitrary scheduling policy. We derive these bounds by analyzing the Increasing Speed Queue (ISQ-$k$). By the WINE formula (Proposition 10.1), this yields the first non-trivial lower bounds on mean response time of the $M/G/k$ system. We start by lower bounding the 2-server system before moving on to the $k$-server system. For the 2-server system, these lower bounds on relevant work are as follows:
@@
We have the following lower bounds of $\mathbb{E}[W_{x}^{M/G/2/OPT}]$:
$E[W_{x}^{M/G/2/OPT}]$ $\geq\dfrac{\lambda_{x}\mathbb{E}[S_{x}^{2}]}{2(1-\rho_{x})}+\dfrac{\mathbb{E}[S_{x}]-(1-\widetilde{S_{x}}(2\lambda_{x}))/2\lambda_{x}}{3-\widetilde{S_{x}}(2\lambda_{x})}+(\lambda-\lambda_{x})x^{2}$ (4.1)
$\mathbb{E}[W_{x}^{M/G/2/OPT}]$ $\geq\dfrac{\lambda_{x}E[S_{x}^{2}]}{2(1-\rho_{x})}+\dfrac{\mathbb{E}[S_{x}]-(1-\widetilde{S_{x}}(2\lambda_{x}))/2\lambda_{x}}{3-\widetilde{S_{x}}(2\lambda_{x})}\cdot\dfrac{1-\rho_{\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu}}{1-\rho_{x}}+\dfrac{(\lambda-\lambda_{x})x^{2}}{2(1-\rho_{x})}.$ (4.2)
Theorem 4.1 is proved in Section 7. Equations 4.1 and 4.2 are derived by applying the drift method to the Sep-ISQ-2 system and AR-ISQ-2 system, respectively. These ISQ-2 variants were introduced in Section 3.3.
The two lower bounds have different strengths at different relevancy-cutoff levels. For small values of $x$, (4.1) is stronger than (4.2). Because the second term in (4.2) is discounted by $\frac{1-\rho_{\bar{x}}}{1-\rho_{x}}$, when $x$ is small $\rho_{\bar{x}}$ is significantly larger than $\rho_{x}$, lowering the numerator. When $x$ is large, this fraction $\frac{1-\rho_{\bar{x}}}{1-\rho_{x}}$ will converge to 1. On the other hand, when $x$ is large, the last term of (4.2) will scale by $\frac{1}{2(1-\rho_{x})}$, whereas this scaling factor is absent in (4.1).
@@
We have the following lower bounds of $\mathbb{E}[W_{x}^{M/G/k/OPT}]$ with $k\geq 3$,
$\mathbb{E}[W_{x}^{M/G/k/OPT}]$ $\geq\dfrac{\lambda_{x}\mathbb{E}[S_{x}^{2}]}{2(1-\rho_{x})}+\dfrac{\lambda_{x}\mathbb{E}[v_{1}(S_{x})]}{2+2\lambda_{x}\mathbb{E}[u_{1}(S_{x})]}+\dfrac{k(\lambda-\lambda_{x})x^{2}}{2},$ (4.3)
$\mathbb{E}[W_{x}^{M/G/k/OPT}]$ $\geq\dfrac{\lambda_{x}\mathbb{E}[S_{x}^{2}]}{2(1-\rho_{x})}+\dfrac{\lambda_{x}\mathbb{E}[v_{1}(S_{x})]}{2+2\lambda_{x}\mathbb{E}[u_{1}(S_{x})]}\cdot\dfrac{1-\rho_{\bar{x}}}{1-\rho_{x}}+\frac{(\lambda-\lambda_{x})J_{x}}{2(1-\rho_{x})},$ (4.4)
@@
In the above theorem, $u_{1}$ and $v_{1}$ is defined by the following recursive formulas,
[…]
with the initial condition $u_{k}(w)=0$ and $v_{k}(w)=0$. The function $h_{k,x}$ is defined in Definition 9.5. Equation 4.3 is based on the Sep-ISQ-$k$ system and Equation 4.4 is based on the AR-ISQ-$k$ system.
Note that $u_{1},v_{1},$ and $h_{k,x}$ can all be exactly symbolically derived for any number of servers, see Appendix C for $u_{1}$ and $v_{1}$ under the 3-server case. See Section 10 for how to convert lower bounds on mean relevant work into lower bounds on mean response time.
4.1. Proof overview
We derive our main results, the lower bounds given in Theorems 4.1 and 4.2, via two steps: First, we prove that the expected relevant work of each ISQ-$k$ system lower bounds the optimal expected relevant work in the $M/G/k$ under any scheduling policy. Second, we characterize the work in the ISQ-$k$ system by applying our novel DiffeDrift method. We now overview the proof in more detail.

### v3

@@ 2.4. Basic adjoint relationship (BAR) and drift method
The well known Basic Adjoint Relationship (BAR) equation states that the stationary distribution $\pi$ of a continuous-time Markov chain $\{Z_{t}\}_{t\geq 0}$ with instantaneous generator $G$ satisfies
@@
under suitable conditions on the Markov chain $\{Z_{t}\}_{t\geq 0}$ and the function $g$ (Glynn2008). By carefully designing the test function $g$, one can solve the BAR equation (2.1) to obtain either exact expressions for the moments of stationary variables or (asymptotically tight) bounds on these moments, depending on the system. This method is commonly referred to as the drift method (Eryilmaz2012, Grosof2023TRaM, Hong2023 and Maguluri2016). We use the drift method to obtain the exact expected total work of the ISQ-$k$ system, and exact results and bounds for further systems we introduce in Section 3. We use these exact results and bounds to lower bound mean response time in the $M/G/k$ under an arbitrary scheduling policy.
@@ 2.4.1. Choosing test functions
Recent applications of the drift method use simple exponential test functions of the form $e^{tQ}$, where $Q$ is the queue length, or $e^{tW}$, where $W$ is the total work in the system. For example, Braverman2017 and Braverman2024 use these test functions to prove heavy-traffic and steady-state approximations in both single-class and multi-class queueing systems. The use of exponential test functions is also known as the transform method (Hurtado-Lange2020 and jhunjhunwala2023).
However, these methods cannot be applied directly to queues with a variable work completion rate, since the test functions $e^{tQ}$ and $e^{tW}$ only yield useful information when the work completion rate is constant. Other recent works have focused on test functions of the form $Q^{2}$ or $W^{2}$ (Grosof2022MSJ), which likewise depend on a constant rate of work completion. Some attempts have been made to circumvent this problem by focusing on heavy traffic settings. For instance, Hurtado-Lange2020 rely heavily on state-space collapse to ensure a constant work completion rate except when the system is near-empty. As the system is rarely empty in heavy traffic, this state-space collapse allows the same straightforward test functions to be used, and heavy traffic results to be obtained.
@@ 2.4.2. Sufficient conditions for BAR
Proposition 3 of Glynn2008 demonstrates a set of sufficient conditions for the BAR equation (2.1) to hold. This proposition has been widely applied in many recent papers in queueing theory. Their result requires the state space to be discrete and certain regularity conditions to hold on the rate matrix of the Markov jump process and the test function $g$. For example, Grosof2023TRaM and Hong2023 invoke Proposition 3 of Glynn2008 by showing that their respective Markov chains have uniformly bounded transition rates. However, this result does not directly apply to our system because our system has a continuous state space.
Theorem 2 of Glynn2008 provides another set of sufficient conditions for the BAR equation (2.1) to hold. This result allows continuous state spaces but requires bounded-drift test functions. However, it is not directly applicable to our unbounded quadratic test function, which has unbounded drift. For test functions with unbounded drift, a standard approach in the existing literature is to truncate these functions, thereby producing a sequence of bounded test functions that approximate the original (Braverman2017 and Guang2024).
@@ 3.1. Queueing model and notation
We say a job is relevant at some threshold $x$ if its remaining size is below $x$. This may occur if the job arrives into the system with an initial size <x<x, or if a job with an initial size >x>x reaches a remaining size below $x$. We refer to the process of a job receiving service and lowering its remaining size as “aging”.
We use $W$ to denote the total work in the system, namely the sum of the remaining sizes of all jobs in the system. We use $W_{x}$ to denote the total remaining size of all relevant jobs in the system, this is the SRPT relevant work. However, our analysis applies to any arbitrary scheduling policy, not just those based on SRPT.
Let $\lambda_{x}:=\lambda F_{S}(x)$ denote the arrival rate of jobs which are relevant from the moment they enter the system. Let $S_{x}:=[S\mid S\leq x]$ denote the conditional job size random variable, truncated at a threshold $x$. The conditional load, $\rho_{x}$, at a threshold $x$ is the average rate at which works with initial size <x<x arrives into the system is given by $\rho_{x}:=\lambda_{x}\mathbb{E}[S_{x}]=\lambda\mathbb{E}[S\mathbbm{1}_{\{S\leq x\}}]$.
We also define the capped job size random variable, $S_{\bar{x}}:=[\min\{S,x\}]$, which represents the amount of relevant work at some threshold $x$ that a random job will contribute over its time in the system. The capped load, $\rho_{\bar{x}}$, is similarly given by $\rho_{\bar{x}}=\lambda\mathbb{E}[S_{\bar{x}}]=\lambda\mathbb{E}[\min\{S,x\}]$.
@@ 3.2. Increasing speed queue
3.3. Auxiliary ISQ systems and recycling
@@
Recall that we call a job relevant if its remaining size is less than some threshold $x$. There are two ways for a job to be relevant: it either enters the system with size less than $x$, or it ages down from an initial size larger than $x$ to a remaining size less than $x$. We refer to the event in which a job with initial size exceeding $x$ ages down to a remaining size below $x$ as recycling, and we call any such job a recycled job. This is standard terminology and see Section 2.3 for how prior work uses the recycling terminology.
@@
We define two auxiliary systems to help us lower bound the response time of the $M/G/k$ system: the separate ISQ-$k$ and the recycling ISQ-$k$.
@@
A separate ISQ-$k$ system (Sep-ISQ-$k$) consists of two subsystems, the truncated-ISQ-$k$ subsystem and the $M/G/\infty$ subsystem. Jobs arrive to the overall Sep-ISQ-$k$ system with the same arrival process as the $M/G/k$, jobs arriving into the system according to a Poisson process with rate $\lambda$ and with job size $S$. Jobs with initial sizes $\leq x$ are routed to an ISQ-$k$ subsystem, which we call the truncated-ISQ-$k$ subsystem. Specifically, the truncated-ISQ-$k$ system is an ISQ-$k$ system characterized by an arrival rate $\lambda_{x}=\lambda F_{S}(x)$, and job size $S_{x}\sim[S\mid S\leq x]$.
@@
On the other hand, jobs with initial sizes >x>x are routed to the separate $M/G/\infty$ subsystem, each server operating at a speed of $1/k$. That is, jobs which are not relevant on arrival are served at a separate subsystem.
@@
The recycling ISQ-$k$ system (Rec-ISQ-$k$) is an ISQ-$k$ system with two arrival streams, a Poisson arrival process with a rate of $\lambda F_{S}(x)$ and job size $S_{x}\sim[S\mid S\leq x]$, and a general arrival process $\{R_{t}\}$ with rate $\lambda(1-F_{S}(x))$ and job sizes exactly $x$. The general arrival process is governed by some general Markovian process $\{R_{t}\}$ for which we only require stability and can be causally dependent on the ISQ-$k$ system, we denote the state of this process as $r$. The idea behind this general arrival process is that it models the moments at which jobs recycle in the M/G/$k$ system under a general Markovian scheduling policy. We refer to the Poisson arrivals as the truncated stream and the general arrival process as the recycling stream.
@@
We will use the triplet $(w,i,r)$ to denote a state of the recycling system where $w$ and $i$ are defined similarly to the states of the full ISQ-$k$ system and $r$ denotes a separate Markovian state from the recycling stream’s state space, which incorporates additional information about the recycling stream. As before, $R$ denotes the corresponding stationary random variable.
@@
In Theorem 6.2, we show that the mean relevant work in the Sep-ISQ-$k$ system lower bounds the mean relevant work in the $M/G/k$ system under any arbitrary scheduling policy, for any arrival rate $\lambda$, job size $S$, and relevancy cutoff $x$. In Theorem 6.3, we prove there always exists a Rec-ISQ-$k$ system lower bounding the mean relevant work of a $M/G/k$ system under an arbitrary scheduling policy. Therefore, our lower bound on the mean relevant work of Rec-ISQ-$k$ system with any recycling process $R$ must lower bound the mean relevant work of any $M/G/k$ system.
@@ 4. Main Results
In this paper, we give the first nontrivial lower bounds on the mean response time of the $M/G/k$ system under an arbitrary scheduling policy. Our bounds are based on lower bounds on relevant work, which we can turn into bounds on response time via the WINE formula (Proposition 5.1). Directly applying this gives the first novel lower bound, the MixEx bound. We improve upon these by proving stronger relevant work lower bounds with the ISQ-$k$ system, these give even stronger bounds, the ISQ and ISQ-Recycling bounds.
We start by lower bounding the 2-server system before moving on to the $k$-server system. For the 2-server system, these lower bounds on mean response time are as follows:
@@
The mean response time of the optimal $M/G/2$ system is lower bounded as follows, which we refer to as the MixEx, ISQ, and ISQ-Recycling bounds, respectively:
$E[T^{M/G/2/OPT}]$ $\geq\dfrac{1}{\lambda}\int_{0}^{\infty}\dfrac{\max\{\mathbb{E}[W_{x}^{M/G/1\text{-SRPT}}],\mathbb{E}[W_{x}^{M/G/\infty}]\}}{x^{2}}\,dx,$ (4.1)
$E[T^{M/G/2/OPT}]$ $\geq\dfrac{1}{\lambda}\int_{0}^{\infty}\dfrac{\max\{\mathbb{E}[W_{x}^{M/G/1\text{-SRPT}}],\mathbb{E}[W_{x}^{M/G/\infty}],\mathbb{E}[W_{x}^{\text{Sep-ISQ-}2}]\}}{x^{2}}\,dx,$ (4.2)
$E[T^{M/G/2/OPT}]$ $\geq\dfrac{1}{\lambda}\int_{0}^{\infty}\dfrac{\max\{\mathbb{E}[W_{x}^{M/G/1\text{-SRPT}}],\mathbb{E}[W_{x}^{M/G/\infty}],\mathbb{E}[W_{x}^{\text{Sep-ISQ-}2}],\mathbb{E}[W_{x}^{\text{Rec-ISQ-}2\text{-L}}]\}}{x^{2}}\,dx,$ (4.3)
where $\mathbb{E}[W_{x}^{M/G/1\text{-SRPT}}]$ and $\mathbb{E}[W_{x}^{M/G/\infty}]$ are standard results given in Section 5 and
$\mathbb{E}[W_{x}^{{\text{Sep-ISQ-}2}}]$ $=\dfrac{\lambda_{x}\mathbb{E}[S_{x}^{2}]}{2(1-\rho_{x})}+\dfrac{\mathbb{E}[S_{x}]-(1-\widetilde{S_{x}}(2\lambda_{x}))/2\lambda_{x}}{3-\widetilde{S_{x}}(2\lambda_{x})}+(\lambda-\lambda_{x})x^{2},$ (4.4)
$\mathbb{E}[W_{x}^{\text{Rec-ISQ-}2\text{-L}}]$ $=\dfrac{\lambda_{x}\mathbb{E}[S_{x}^{2}]}{2(1-\rho_{x})}+\dfrac{\mathbb{E}[S_{x}]-(1-\widetilde{S_{x}}(2\lambda_{x}))/2\lambda_{x}}{3-\widetilde{S_{x}}(2\lambda_{x})}\cdot\dfrac{1-\rho_{\mkern 1.5mu\overline{\mkern-1.5mux\mkern-1.5mu}\mkern 1.5mu}}{1-\rho_{x}}+\dfrac{(\lambda-\lambda_{x})x^{2}}{2(1-\rho_{x})}.$ (4.5)
Proof deferred to Section 8.1.
We call Equation 4.1 the MixEx bound, as it combines two existing lower bounds on mean relevant work, namely the $M/G/1$-SRPT and $M/G/\infty$ relevant work, through the WINE formula. These existing lower bounds, together with the WINE formula, are explained in greater detail in Section 5.
The remaining parts of Theorem 4.1 are proved in Section 8.1, see Section 4.1 for proof outline. We refer to Equation 4.2 as the ISQ bound and to Equation 4.3 as the ISQ-Recycling bound, because each introduces a new lower bound on the mean relevant work obtained from the Sep-ISQ-2 and Rec-ISQ-2 systems, respectively. Specifically, we derive Equations 4.4 and 4.5 by applying our novel variant of the drift method, called the DiffeDrift method, to the Sep-ISQ-2 and Rec-ISQ-2 systems, respectively. These ISQ-2 variants are introduced in Section 3.3. Numerical studies of these bounds are given in Section 12.
@@
The mean response time of the optimal $M/G/k$ system with $k\geq 3$ is lower bounded as follows, which we refer to as the MixEx, ISQ, and ISQ-Recycling bounds, respectively:
$E[T^{M/G/k/OPT}]$ $\geq\dfrac{1}{\lambda}\int_{0}^{\infty}\dfrac{\max\{\mathbb{E}[W_{x}^{M/G/1\text{-SRPT}}],\mathbb{E}[W_{x}^{M/G/\infty}]\}}{x^{2}}\,dx,$ (4.6)
$E[T^{M/G/k/OPT}]$ $\geq\dfrac{1}{\lambda}\int_{0}^{\infty}\dfrac{\max\{\mathbb{E}[W_{x}^{M/G/1\text{-SRPT}}],\mathbb{E}[W_{x}^{M/G/\infty}],\mathbb{E}[W_{x}^{\text{Sep-ISQ-}k}]\}}{x^{2}}\,dx,$ (4.7)
$E[T^{M/G/k/OPT}]$ $\geq\dfrac{1}{\lambda}\int_{0}^{\infty}\dfrac{\max\{\mathbb{E}[W_{x}^{M/G/1\text{-SRPT}}],\mathbb{E}[W_{x}^{M/G/\infty}],\mathbb{E}[W_{x}^{\text{Sep-ISQ-}k}],\mathbb{E}[W_{x}^{\text{Rec-ISQ-}k\text{-L}}]\}}{x^{2}}\,dx,$ (4.8)
where $\mathbb{E}[W_{x}^{M/G/1\text{-SRPT}}]$ and $\mathbb{E}[W_{x}^{M/G/\infty}]$ are standard results given in Section 5 and
$\mathbb{E}[W_{x}^{\text{Sep-ISQ-}k}]$ $=\dfrac{\lambda_{x}\mathbb{E}[S_{x}^{2}]}{2(1-\rho_{x})}+\dfrac{\lambda_{x}\mathbb{E}[v_{1}(S_{x})]}{2+2\lambda_{x}\mathbb{E}[u_{1}(S_{x})]}+\dfrac{k(\lambda-\lambda_{x})x^{2}}{2},$ (4.9)
$\mathbb{E}[W_{x}^{\text{Rec-ISQ-}k\text{-L}}]$ $=\dfrac{\lambda_{x}\mathbb{E}[S_{x}^{2}]}{2(1-\rho_{x})}+\dfrac{\lambda_{x}\mathbb{E}[v_{1}(S_{x})]}{2+2\lambda_{x}\mathbb{E}[u_{1}(S_{x})]}\cdot\dfrac{1-\rho_{\bar{x}}}{1-\rho_{x}}+\frac{(\lambda-\lambda_{x})J_{x}}{2(1-\rho_{x})},$ (4.10)
@@
Proof deferred to Section 10.1.
In the above theorem, we define $u_{1}$ and $v_{1}$ via the following recursive formulas,
$u_{q}(w)$ $=e^{-\frac{kw\lambda_{x}}{q}}\int_{0}^{w}\dfrac{e^{\frac{k\lambda_{x}y}{q}}(k-q+k\lambda_{x}\mathbb{E}[u_{q+1}(S_{x}+y)])}{q}\,dy,$
$v_{q}(w)$ $=e^{-\frac{kw\lambda_{x}}{q}}\int_{0}^{w}\dfrac{e^{\frac{k\lambda_{x}y}{q}}(2ky-2qy+k\lambda_{x}\mathbb{E}[v_{q+1}(S_{x}+y)])}{q}\,dy,$
with the initial condition $u_{k}(w)=0$ and $v_{k}(w)=0$ (see Definitions 10.1 and 10.2). Both $u_{1}$ and $v_{1}$ are also derived using the DiffeDrift method. The function $h_{k,x}$ is defined in Definition 10.5. Equation 4.9 is based on the Sep-ISQ-$k$ system and Equation 4.10 is based on the Rec-ISQ-$k$ system.
Note that $u_{1},v_{1},$ and $h_{k,x}$ can all be exactly symbolically derived for any number of servers, see Appendix C for $u_{1}$ and $v_{1}$ under the 3,4 and 5-server cases. See Section 5 for how to convert lower bounds on mean relevant work into lower bounds on mean response time. In Section 11, we further discuss how a specialized program can be developed to compute these recursive formulas efficiently for $k\geq 6$, and how our framework extends beyond the known-size setting.
4.1. Proof outline
To derive lower bounds on mean response time, we first obtain lower bounds on the mean relevant work. By the WINE formula (see Section 5), these directly translate into lower bounds on mean response time. Applying this framework to the naive lower bounds (Equations 5.2 and 5.3) already yields some improvement in certain load regimes. Our approach, however, goes further, providing substantially tighter lower bounds.
We derive the novel relevant work lower bounds given in Theorems 4.1 and 4.2, via two steps: First, we prove that the expected relevant work of each ISQ-$k$ system lower bounds the optimal expected relevant work in the $M/G/k$ under any scheduling policy. Second, we characterize the work in the ISQ-$k$ system by applying our novel DiffeDrift method.